OpenStudy (anonymous):

29w + 33 = -14w + 162 W=? @Shay17

OpenStudy (anonymous):

w=129/43

OpenStudy (anonymous):

Yes, the answer is 3.

OpenStudy (anonymous):

29w+33=-14w+162 subtract 33 29w=-14w+129 add14w 43w=129 divide 43 w=3